# Commutation relations for two-sided radial SLE

### v2

@@ 4.2 Semiclassical limits of partition functions
Then the result follows from the large deviation principle for chordal SLE as $\kappa\to 0{\scriptstyle+}$ [29]. In fact, the Loewner energy is the large deviation rate function of chordal $\operatorname{SLE}_{0{\scriptstyle+}}$ for the Hausdorff metric and $\gamma\mapsto-\log\mathrm{CR}(\mathbb{D}\setminus\gamma)$ is a continuous function $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})\to[0,\infty]$. Varadhan’s lemma [8, Lem. 4.3.4 and 4.3.6] shows if $\alpha\sim-\lambda/\kappa$, then
@@
and has the left derivative $-\sqrt{2\lambda}$ and right derivative $\sqrt{2\lambda}$ at $\theta=\pi$.
@@
If one removes the assumption of interchangeability from Proposition 4.1, then one obtains, in addition to the solutions (4.2) and (4.3), the solutions
[…]
(These solutions correspond to Eq. (4.10) with $C>-3$ and $\lambda=(3+C)/2$.) Note that $\mathcal{U}^{\lambda,+}(\theta_{1},\theta_{2})=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$ for $\theta_{1}<\theta_{2}\leq\theta_{1}+\pi$ and $\mathcal{U}^{\lambda,-}(\theta_{1},\theta_{2})=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$ for $\theta_{1}+\pi\leq\theta_{2}\leq\theta_{1}+2\pi$. As in the case with interchangeability, we do not have a one-to-one correspondence with the solutions from the $\kappa>0$ case. Following Lemma 4.7, one can take the semiclassical limit of $\mathcal{Z}_{\alpha,\beta}$ from Proposition 3.17. If $\lambda>0$ and $\alpha\sim-\lambda/\kappa$ as $\kappa\to 0$, then one has
$\lim_{\kappa\to 0}\kappa\log\frac{\mathcal{Z}_{\alpha,\beta}(\theta_{1},\theta_{2})}{\mathcal{Z}_{\alpha,\beta}(0,\pi)}=\begin{cases}\mathcal{U}^{\lambda,+}(\theta_{1},\theta_{2}),\quad&\text{if }\beta=1;\\ \mathcal{U}^{\lambda}(\theta_{1},\theta_{2}),\quad&\text{if }\beta\in(0,1);\\ \mathcal{U}^{\lambda,-}(\theta_{1},\theta_{2}),\quad&\text{if }\beta=0.\end{cases}$
The case $\beta=1$ follows from
@@
which can be seen by examining the proof of Proposition 4.8. The case $\beta=0$ is analogous.
@@ Appendix A Euler’s hypergeometric differential equations
$-\delta:=f^{\prime}(1/2+\varepsilon)<0.$
@@
Appendix B Convergence of radial SLE to chordal SLE
Suppose $\gamma:[0,\infty)\to{\mathbb{H}}$ is a continuous non-self-crossing curve in ${\mathbb{H}}$ such that $\gamma_{0}=0$. Let $H_{t}$ be the unbounded connected component of ${\mathbb{H}}\setminus\gamma_{[0,t]}$. Let $g_{t}:H_{t}\to{\mathbb{H}}$ be the unique conformal map with $\lim_{z\to\infty}|g_{t}(z)-z|=0$. We say that the curve is parameterized by half-plane capacity if
$g_{t}(z)=z+\frac{2t}{z}+o\left(\frac{1}{|z|}\right),\quad\text{as }z\to\infty.$
Then $g_{t}$ satisfies the chordal Loewner equation
$\partial_{t}g_{t}(z)=\frac{2}{g_{t}(z)-W_{t}},\quad g_{0}(z)=z.$
We call $W_{t}$ the driving function. Chordal $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $\infty$ is the chordal Loewner chain with driving function $W_{t}=\sqrt{\kappa}B_{t}$ where $B$ is one-dimensional Brownian motion. For a general simply connected domain $D$ with two distinct prime ends $x_{1},x_{2}\in\partial D$, we define chordal $\operatorname{SLE}_{\kappa}$ in $D$ from $x_{1}$ to $x_{2}$ to be the image of chordal $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $\infty$ under the conformal map $\varphi:{\mathbb{H}}\to D$ sending 00 to $x_{1}$ and $\infty$ to $x_{2}$.
Proof of Lemma 3.11.
Recall that $Q_{t_{1}}$ denotes the law of radial $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]};\mathrm{e}^{\mathfrak{i}\theta_{2}};0)$, and denote by $Q_{\infty}$ the law of chordal $\operatorname{SLE}_{\kappa}$ in $(\mathbb{D}\setminus\eta^{(1)};\mathrm{e}^{\mathfrak{i}\theta_{2}},0)$. We fix a sequence of conformal maps $f_{t_{1}}:\mathbb{D}\setminus\eta^{(1)}_{[0,t_{1}]}\to{\mathbb{H}}$ and a conformal map $f_{\infty}:\mathbb{D}\setminus\eta^{(1)}\to{\mathbb{H}}$ such that
$f_{t_{1}}\left(\eta^{(1)}_{t_{1}}\right)=-1,\quad f_{t_{1}}(\mathrm{e}^{\mathfrak{i}\theta_{2}})=0;\qquad f_{\infty}(0)=-1,\quad f_{\infty}(\mathrm{e}^{\mathfrak{i}\theta_{2}})=0;$
and that $f_{t_{1}}$ converges to $f_{\infty}$ locally uniformly. Note that $(f_{t_{1}})_{*}(Q_{t_{1}})$ is the same as radial $\operatorname{SLE}_{\kappa}$ in $({\mathbb{H}};0;w=f_{t_{1}}(0))$, and $(f_{\infty})_{*}(Q_{\infty})$ is the same as chordal $\operatorname{SLE}_{\kappa}$ in $({\mathbb{H}};0,-1)$.
Suppose $\gamma$ is chordal $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $\infty$ and denote by $W_{t}$ its driving function and by $g_{t}$ the corresponding conformal maps. From [36], the law of chordal $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $-1$ is the same as $\gamma$ weighted by the local martingale
$|g_{t}^{\prime}(-1)|^{h}(W_{t}-g_{t}(-1))^{-2h},$ (B.1)
where $h=(6-\kappa)/(2\kappa)$ as in (3.1). The law of radial $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $w=f_{t_{1}}(0)$ is the same as $\gamma$ weighted by the local martingale
$|g_{t}^{\prime}(w)|^{\tilde{h}}\times\left(\frac{\operatorname{Im}(g_{t}(w))}{\operatorname{Im}(w)}\right)^{(\kappa-6)^{2}/(8\kappa)}\times\left|\frac{W_{t}-\operatorname{Re}(g_{t}(w))}{\operatorname{Re}(w)}\right|^{-2h},$ (B.2)
where $\tilde{h}=(\kappa-2)(6-\kappa)/(8\kappa)$ as in (3.1).
Combining (B.1) and (B.2), $(f_{t_{1}})_{*}(Q_{t_{1}})$ (radial $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $w=f_{t_{1}}(0)$) is the same as $(f_{\infty})_{*}(Q_{\infty})$ (chordal $\operatorname{SLE}_{\kappa}$ in ${\mathbb{H}}$ from 00 to $-1$) weighted by
$M_{t}=\frac{|g_{t}^{\prime}(w)|^{\tilde{h}}}{|g_{t}^{\prime}(-1)|^{h}}\times\left(\frac{\operatorname{Im}(g_{t}(w))}{\operatorname{Im}(w)}\right)^{(\kappa-6)^{2}/(8\kappa)}\times\left|\frac{W_{t}-\operatorname{Re}(g_{t}(w))}{(W_{t}-g_{t}(-1))\operatorname{Re}(w)}\right|^{-2h}.$
As $t_{1}\to\infty$, we have $w=f_{t_{1}}(0)\to-1$ and $M_{t}\to 1$ almost surely, and $f_{t_{1}}$ converges to $f_{\infty}$ locally uniformly. These give the desired conclusion. ∎
@@ Acknowledgments.
  • We thank Yu Feng, Chongzhi Huang, Eveliina Peltola, Fredrik Viklund, and Lu Yang for helpful discussions. We thank Minjae Park for the simulation in Figure 2.
@@
  • H.W. is supported by Beijing Natural Science Foundation (JQ20001). Part of this work was carried out during H.W.’s visit at IHES under the support of “K.C. Wong Education Foundation”. H.W. is partly affiliated with Yanqi Lake Beijing Institute of Mathematical Sciences and Applications, Beijing, China.

### v3

@@ 4.2 Semiclassical limits of partition functions
Then the result follows from the large deviation principle for chordal SLE as $\kappa\to 0{\scriptstyle+}$ [34]. In fact, the Loewner energy is the large deviation rate function of chordal $\operatorname{SLE}_{0{\scriptstyle+}}$ for the Hausdorff metric and $\gamma\mapsto-\log\mathrm{CR}(\mathbb{D}\setminus\gamma)$ is a continuous function $\mathfrak{X}(\mathbb{D};\mathrm{e}^{\mathfrak{i}\theta_{1}},\mathrm{e}^{\mathfrak{i}\theta_{2}})\to[0,\infty]$. Varadhan’s lemma [11, Lem. 4.3.4 and 4.3.6] shows if $\alpha\sim-\lambda/\kappa$, then
@@
and has the left derivative $\sqrt{2\lambda}$ and right derivative $\sqrt{2\lambda}$ at $\theta=\pi$.
@@
Recall from Remark 4.2 that if one removes the assumption of interchangeability from Proposition 4.1, then one obtains, in addition to the solutions (4.2) and (4.3), the solutions $\mathcal{U}_{\lambda,\pm}$, $\lambda>0$. Note that $\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2})=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$ for $\theta_{1}<\theta_{2}\leq\theta_{1}+\pi$ and $\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2})=\mathcal{U}^{\lambda}(\theta_{1},\theta_{2})$ for $\theta_{1}+\pi\leq\theta_{2}<\theta_{1}+2\pi$. Following Lemma 4.8, one can take the semiclassical limit of $\mathcal{Z}_{\alpha,\beta}$ from Proposition 3.16. For example, if $\lambda>0$ and $\alpha\sim-\lambda/\kappa$, we have
$\lim_{\kappa\to 0}\kappa\log\frac{\mathcal{Z}_{\alpha,1}(\theta_{1},\theta_{2})}{\mathcal{Z}_{\alpha,1}(0,\pi)}=\mathcal{U}_{\lambda,+}(\theta_{1},\theta_{2}).$
This follows from
@@
which can be seen by examining the proof of Proposition 4.9. By a similar analysis one finds
$\lim_{\kappa\to 0}\kappa\log\frac{\mathcal{Z}_{\alpha,0}(\theta_{1},\theta_{2})}{\mathcal{Z}_{\alpha,0}(0,\pi)}=\mathcal{U}_{\lambda,-}(\theta_{1},\theta_{2}).$
@@ Appendix A Euler’s hypergeometric differential equations
$\delta:=f^{\prime}(1/2+\varepsilon)<0.$
@@
Appendix B Solutions of the separable equation (4.7)
Lemma B.1.
Let $C\in\mathbb{R}$ and consider the separable differential equation
$(8+2C\sin^{2}(\theta/2))f^{\prime}(\theta)=(f(\theta))^{2}+C+4.$ (B.1)
The only $C^{1}$ solutions of (B.1) defined on $(0,2\pi)$ are
  • if $C+4>0$,
[…]
    for any $D\in[-\pi/4,\pi/4],$
  • if $C+4=0$,
[…]
    for any $D\leq 0$ and $E\geq 0$,
  • if $C+4<0$,
    $f(\theta)=\sqrt{-C-4}\quad\text{and}\quad f(\theta)=-\sqrt{-C-4}.$
Proof.
Note that
$\frac{1}{8+2C\sin^{2}(\theta)}=\frac{1}{4+(C+4)\tan^{2}(\theta/2)}\frac{\mathrm{d}}{\mathrm{d}\theta}\tan(\theta/2).$
Thus, we can then rewrite (B.1) as
$\frac{\mathrm{d}f}{f^{2}+(C+4)}=\frac{\mathrm{d}\tan(\theta/2)}{4+(C+4)\tan^{2}(\theta/2)},$ (B.3)
under the assumptions $(f(\theta))^{2}\neq-(C+4),$ $\theta\neq\pi$ and $(C+4)\tan^{2}(\theta/2)\neq-4.$
Case 1: Suppose $C+4>0$. By the Picard-Lindelöf theorem Eq. (4.7) has a unique local solution $f$ around $\theta_{0}\in(0,2\pi)$ for each initial condition $f(\theta_{0})=f_{0}$. Integrating both sides of (B.3) gives
$\arctan(f(\theta)/\sqrt{C+4})=\begin{cases}\frac{\arctan(\frac{1}{2}\sqrt{C+4}\tan(\theta/2))}{2}+D-\pi/4&\theta\in(0,\pi)\\ \frac{\arctan(\frac{1}{2}\sqrt{C+4}\tan(\theta/2))}{2}+D+\pi/4&\theta\in(\pi,2\pi)\end{cases}$
for some $D$. Note that the left-hand side takes values in $(\-\pi/2,\pi/2)$ and that this is true for the right-hand side only if $D\in[-\pi/4,\pi/4]$. Thus the only global solutions of (B.1) in the case $C+4>0$ are the ones given in (B.2).
Case 2: Suppose $C+4=0$. Picard-Lindelöf guarantees existence and uniqueness of local solutions to the initial value problem away from $\theta=\pi$. The solutions on $(0,\pi)$ are of the form $f\equiv 0$ or
$f(\theta)=\frac{1}{D-\frac{1}{4}\tan(\theta/2)}$
for some $D\in\mathbb{R}$. In order for the latter to be defined on all of $(0,\pi)$ we require $D\leq 0$. Similarly, the solutions on $(\pi,2\pi)$ are of the form $f\equiv 0$ or
$f(\theta)=\frac{1}{E-\frac{1}{4}\tan(\theta/2)}$
for some $E\geq 0$. To find solutions on $(0,2\pi)$ we patch the solutions above. For this we note that the left and right derivatives of $f(\theta)=(D-\frac{1}{4}\tan(\theta/2))^{-1}$ at $\pi$ are both $2$, regardless of $D$.
Case 3: Finally, fix $C+4<0$ and let $\theta_{0}$ be the unique solution to $8+2C\sin^{2}(\theta/2)=0$ on $(0,\pi)$. Then Picard-Lindelöf guarantees existence and uniqueness of local solutions to the initial value problem around each $\theta\in(0,2\pi)\setminus\{\theta_{0},2\pi-\theta_{0}\}$. For the initial values $\pm\sqrt{-C-4}$ we have constant solutions. For the other initial values we integrate (B.3) and find solutions of the form
$\log\frac{|f(\theta)-\sqrt{-C-4}|}{|f(\theta)+\sqrt{-C-4}|}=\frac{1}{2}\log\frac{|\sqrt{-C-4}\tan(\theta/2)+2|}{|\sqrt{-C-4}\tan(\theta/2)-2|}+D,$
$D\in\mathbb{R}$. More concretely, if $\theta_{1}\in(0,2\pi)\setminus\{\theta_{0},2\pi-\theta_{0}\}$ and $f(\theta_{1})\in\mathbb{R}\setminus\{\pm\sqrt{-C-4}\}$ then
$f(\theta)=T\bigg(\frac{1}{2}\log\frac{|\sqrt{-C-4}\tan(\theta/2)+2|}{|\sqrt{-C-4}\tan(\theta/2)-2|}+D\bigg)$
for $\theta$ in the connected component of $(0,2\pi)\setminus\{\theta_{0},2\pi-\theta_{0}\}$ containing $\theta_{1}$, where $T$ is the inverse of $\log\frac{|x-\sqrt{-C-4}|}{|x+\sqrt{-C-4}|}$ on the connected component of $\mathbb{R}\setminus\{\pm\sqrt{-C-4}\}$ containing $f(\theta_{1})$. An explicit computation shows that all such solutions satisfy $|f^{\prime}(\theta)|\to\infty$ as $\theta\to\theta_{0}$ or $\theta\to 2\pi-\theta_{0}$ and therefore they can not be extended to $C^{1}$ solutions on $(0,2\pi).$
@@ Acknowledgments.
  • We thank Mo Chen, Yu Feng, Chongzhi Huang, Eveliina Peltola, Fredrik Viklund, and Lu Yang for helpful discussions. We thank Minjae Park for the simulation in Figure 2. We thank the anonymous referees for their constructive comments on a previous version of the paper.
@@
  • H.W. is supported by Beijing Natural Science Foundation (JQ20001) and New Cornerstone Investigator Program 100001127. Part of this work was carried out during H.W.’s visit at IHES under the support of “K.C. Wong Education Foundation”. H.W. is partly affiliated with Yanqi Lake Beijing Institute of Mathematical Sciences and Applications, Beijing, China.
Conflict of interest. The authors have no competing interests to declare that are relevant to the content of this article.